# Rapid Word Learning Through Meta In-Context Learning

### v1

@@ Results
For the quantitative evaluation, we first present the 1-shot scores of comparing the model-generated and ground-truth definitions for Llama-3 8B baselines and the Minnow models in Table 3. In Appendix G, we present 1-shot scores for all models (including the specialized FLAN-T5 models and the original CoLLEGe model) in Table 13 and averaged 1-, 2-, and 3-shot results on CoLLEGe-DefGen in Table 14. In all of these settings, Minnow finetuning improves the baseline scores by 0.3–1.5 on BERTScore F1 and 3.1–5.3 on ROUGE-L. On CoLLEGe-DefGen, the Minnow model finetuned from the instruction-tuned Llama-3 8B outperforms all other models across all settings. On Oxford, the Minnow models finetuned from both variants perform comparably well, but they are inferior to the largest specialized FLAN-T5 by 2.9 on ROUGE-L. However, note that our Minnow finetuning is neither tailored for generating definitions nor using these specific definition datasets. In Table 1, when comparing the definitions generated from each baseline and a Minnow model finetuned from that baseline, the latter is more often favored over the corresponding baselines for both Llama-3 8B variants.
For the qualitative evaluation, Table 4 shows Minnow-model-generated and ground-truth definitions for words from CoLLEGe-DefGen (see Tables 15 and 16 in Appendix G for additional examples from CoLLEGe-DefGen and Oxford). To summarize our manual analysis, we find that definitions generated by the Minnow model often capture most of the word meanings, form reasonable inferences from the contexts, and outperform the baseline. However, they are not always precise compared to the ground-truth definitions.
@@ 6 Conclusion
The efficacy of Minnow, or meta-learning in general, suggests that human-level efficiency in linguistic generalizations may be acquired through practicing over many instances of learning tasks, without presuming strict, explicit inductive biases (Russin et al., 2024; Irie and Lake, 2024). Whether models achieve the generalizations in this work through human-like mechanisms, such as systematicity and categorical abstraction, remains for future analysis.
@@ 7 Limitations
In addition, we only consider learning a single new word on the fly. However, in real-world learning, both humans and models need to continually learn multiple words, usages, and even abstract rules (Mueller et al., 2024). Implementing this continual learning setting would be another future direction.
@@
When testing LLMs (Section 5), the words and example sentences we use may already exist in the pre-training data, potentially allowing LLMs to recall known word meanings rather than learn genuinely new ones (note, however, the Chimera dataset introduces new concepts which are unusual and not lexicalized). The performance of the baseline LLMs shows that, even with this potential worry, there is room for improvement, which the Minnow-finetuned LLMs are able to achieve.
@@ Acknowledgements
We thank Michael Hu, Will Merrill, Sophie Hao, Byung-Doh Oh, Shauli Ravfogel, and other members of the Computation and Psycholinguistics Lab for insightful and helpful discussions and comments. This work is supported by the National Science Foundation under NSF Award 1922658 (for Wentao Wang) and IIS-2239862. This work is also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
@@ Appendix A Word Usage Dataset Creation
Since the basic units of our focus are words (as opposed to word pieces in other tokenization schemes), we need to identify words in the text. To achieve this, we apply the same word-level tokenization to all datasets (for consistency) and mark word boundaries by whitespace during preprocessing. Models trained from scratch use this word-level tokenization. When the text is used in finetuning Llama-3, which comes with its pre-trained subword tokenizer, we remove the unnatural spaces introduced by the word-level tokenization and tokenize the text again with Llama-3 tokenizer, so the text format becomes closer to its pre-training data (See the Finetuning paragraph in Appendix B for further details of this process). For CHILDES data, we preprocess the data in the same way as Yedetore et al. (2023) did, which uses children’s input in the North American English portion, but we do not split and unk the data at the preprocessing stage. For BabyLM data, we use the data in the 10M track of the BabyLM Challenge 2023, which mixes 10 portions, each from a different data source (child- or adult-oriented, speech transcription or written text like Wikipedia). We exclude the QED portion for its poor quality (also mentioned in the 2nd BabyLM Challenge). We apply word-level tokenization on untokenized portions, and then split the text into sentences using heuristics. We use spaCy for all word-level tokenization along with Part-Of-Speech tagging. We lowercase all text before preprocessing to unify the capitalization of words in different places. We deduplicate sentences and remove sentences having less than 1 word (not counting punctuation).
@@ Appendix B Model and Training Configurations
We finetune Llama-3 8B (Llama Team, Meta AI, 2024) with Minnow on each of the CHILDES and BabyLM-10M datasets, but we refer to the models finetuned on BabyLM-10M by default, as we mentioned in Section 5. We finetune from both the pre-trained and instruction-tuned variants of Llama-3 8B, but we refer to the models finetuned from the pre-trained variant by default, presenting results of finetuning from the instruction-tuned variant only in the generative settings, where their performance may differ significantly due to their different capabilities to follow the prompt. We use two reserved special tokens in Llama-3 tokenizer vocabulary as the placeholder token and the separator token. To make the tokenization more natural to the model’s pre-training data, we clean up tokenization spaces in the text (e.g., the space before “,”, “.”, or “’s”) introduced by the word-level tokenization during preprocessing and make the placeholder token absorbs any preceding spaces of the word. Finetuning is minimally parameter-efficient: We finetune only the input and output embeddings of the two special tokens, while freezing all other parameters. Before finetuning, the input/output embedding of either token is initialized to the mean of all input/output embeddings (Hewitt, 2021). When finetuning the model on CHILDES with 5 examples per episode, we use batch size (i.e., number of episodes/sequences per batch) $32$ and initial learning rate $3\times 10^{-3}$ and truncate the sequence to the max length of $80$ tokens to control the memory usage. When finetuning the model on CHILDES with 10 examples per episode, we use batch size $8$ and initial learning rate $3\times 10^{-4}$ and truncate the sequence to the max length of $180$ tokens. When finetuning the model on BabyLM-10M with $5$ examples per episode, we use batch size $16$ and initial learning rate $1\times 10^{-3}$ and truncate the sequence to the max length of $160$ tokens. Other settings are the same as when training from scratch except that we do not apply weight decay. Each run is performed on a single A100 GPU for 15 epochs on CHILDES (33 hours) or 12 epochs on BabyLM-10M (48 hours).
@@ Appendix C Held-out Word Classification
This is the baseline model introduced in Section 3.4. We follow the format described there and additionally prepend a prompt to make the performance better: “The following lines are lowercased example sentences using a new word ‘dax’ in random order, one per line:”. (We discuss the consequence of using a same pseudo-word in Appendix H.)
@@
dataset $K$ $C$ Minnow from scratch Llama-3 8B with ‘’ Llama-3 8B with ‘dax Llama-3 8B +Minnow
CHILDES $5$ 4 72.3(1.6) 58.33 71.09 79.1(0.5)
8 59.8(0.4) 46.49 60.02 70.4(0.2)
$10$ 4 75.1(0.7) 66.56 76.53 84.9(0.2)
8 63.4(1.5) 56.17 66.05 75.9(0.6)
BabyLM-10M $5$ 4 77.4(0.5) 70.45 78.39 86.5(0.6)
8 67.5(0.7) 60.12 69.74 80.5(1.0)
@@ Appendix D Syntactic Category Classification
Figure 2: Syntactic classification accuracy. Error bar shows the 95% confidence interval given 3 runs. “Minnow from scratch on CHILDES” (blue) and “Minnow from scratch on BabyLM-10M” (orange) mean the models trained from scratch with Minnow on CHILDES and BabyLM-10M, respectively. (These models have a closed vocabulary, so many words in the dataset will be Out-Of-Vocabulary and be presented as <unk>, which could make the task easier.) “Llama-3 8B baseline” (green) means Llama-3 8B baseline with pseudo-word “dax”. “Llama-3 8B +Minnow on BabyLM-10M” (red) means Llama-3 8B finetuned with Minnow on BabyLM-10M. “N”, “V”, “Adj”, and “Adv” are short for noun, verb, adjective, and adverb, respectively. “Mean” is the mean across all category pairs. The black dashed line marks the chance level (50%). “Llama-3 8B +Minnow on BabyLM-10M” (red) shows improvement over “Llama-3 8B baseline” (green) in all category pairs, with mean accuracy risen from 64% to 83%. Note that “Minnow from scratch on BabyLM-10M” (orange) has a 77% mean accuracy, much better than the baseline accuracy and even comparable to the Minnow models finetuned from Llama-3 8B on many category pairs, again demonstrating its data efficiency.
@@
Noun Verb 71.0 43 99 86.3(1.5) 74.7(1.7) 98.0(1.6)
Noun Adjective 66.0 79 53 84.0(2.2) 71.3(4.6) 96.7(0.5)
Noun Adverb 64.0 55 73 81.3(2.2) 75.7(1.7) 87.0(2.9)
Verb Adjective 70.5 49 92 92.7(0.5) 90.0(2.2) 95.3(1.2)
Verb Adverb 53.0 85 21 78.8(5.2) 90.0(2.4) 67.7(12.5)
Adjective Adverb 61.5 42 81 72.8(0.2) 57.3(2.6) 88.3(3.1)
Table 7: LLMs’ accuracies (%) of distinguishing two syntactic categories in novel contexts. We show the mean and the standard deviation (in the bracket) of 3 runs. ‘Acc. (1$>$2)’ denotes the accuracy on the set of sentences where Category 1 should be preferred over Category 2 (e.g., assigning a higher probability to a noun in a noun-expecting context for row 1), and vice versa. Column ‘Acc.’ lists the aggregate accuracy. “Llama-3 8B +Minnow” have accuracies significantly better than chance except distinguishing adjective from verb (row 5). Additionally, “Llama-3 8B +Minnow” improves over Llama-3 8B baseline in differentiating most category pairs except discriminating nouns from adjectives (row 2), showing the effectiveness of finetuning with Minnow.
@@
Table 8: Accuracies (%) of distinguishing two syntactic categories in novel contexts for models trained from scratch with Minnow. We show the mean and the standard deviation (in the bracket) of 3 runs. ‘Acc. (1$>$2)’ denotes the accuracy on the set of sentences where Category 1 should be preferred over Category 2 (e.g., assigning higher probability to a noun in a noun-expecting context for row 1), and vice versa. Column ‘Acc.’ lists the aggregate accuracy. Both models perform better than chance on many category pairs, suggesting that models can develop some ability to one-shot learn the syntactic category of a word from human-scale data with Minnow.
@@ Appendix E Comparing Generations
Tables 1 and 9 show the results of comparing Llama-3 8B baseline (pre-trained to instruction-tuned) to the Minnow model finetuned from that baseline (with random seed 0) on new examples and definitions generated by greedy decoding and top-p=$0.92$, respectively.
@@
Instruction-tuned baseline 46 52 33

### v2

@@ Results
For the quantitative evaluation, we first present the 1-shot scores of comparing the model-generated and ground-truth definitions for Llama-3 8B and CoLLEGe baselines, the Minnow models, and one specialized model in Table 3. In Appendix G, we present 1-shot scores for all models (Table 16) and averaged 1-, 2-, and 3-shot results on CoLLEGe-DefGen (Table 17). Minnow finetuning improves the Llama-3 8B baseline by 0.3–1.5 on BERTScore F1 and 3.1–5.3 on ROUGE-L. On CoLLEGe-DefGen, the Minnow model finetuned from the instruction-tuned Llama-3 8B outperforms all other non-specialized models across all settings. On Oxford, the Minnow models finetuned from both variants of Llama-3 8B perform comparably well, but they are inferior to the largest specialized model by 2.9 on ROUGE-L. However, note that our Minnow finetuning is neither tailored for generating definitions nor using these definition datasets. In Table 1, the Minnow model is more often favored over each corresponding baseline.
For the qualitative evaluation, Table 4 shows Minnow-model-generated and ground-truth definitions for a word from CoLLEGe-DefGen (see Tables 18 and 19 in Appendix G for additional examples from CoLLEGe-DefGen and Oxford). In our manual analysis, we find that definitions generated by the Minnow model often capture most of the word meanings, form reasonable inferences from the contexts, and outperform the baselines. However, they are not always precise compared to the ground-truth definitions.
@@ 6 Conclusion
The efficacy of Minnow, or meta-learning in general, suggests that human-level efficiency in linguistic generalizations may be acquired through practicing over many instances of learning tasks, without presuming strict, explicit inductive biases (Russin et al., 2024; Irie and Lake, 2024). Whether models achieve the generalizations in this work through human-like mechanisms remains for future analysis.
@@ 7 Limitations
In addition, we only consider learning a single new word on the fly, and each word is represented by the same special token. However, in real-world learning, both humans and models need to continually learn multiple words, usages, and even abstract rules (Sinha et al., 2023; Lampinen, 2024; Mueller et al., 2024). Future work could implement continual learning of multiple different words by meta-training to learn real words or pseudo-words. This way of meta-training could also allow the rapid word learning ability to co-exist with other general abilities of language models.
@@
When testing LLMs (Section 5), the words and example sentences we use may already exist in the pre-training data, potentially allowing LLMs to recall known word meanings rather than learn genuinely new ones [^10] Eisenschlos et al. (2023) suggested a similar solution called the reverse dictionary (Hill et al., 2016), through which the model may identify the underlying concept from the word definition. (note, however, the Chimera dataset introduces new concepts which are unusual and not lexicalized). The performance of the baseline LLMs shows that, even with this potential worry, there is room for improvement, which the Minnow-finetuned LLMs are able to achieve.
@@
Morphological Features
In the paper, we focus on learning word-forms, each represented by a single special placeholder token, without considering morphological features. However, different words (word-forms) may be derived from the same root word and thus share a common underlying concept, but differ only in their affixes. Such a group of words forms a word family (Bauer and Nation, 1993), and such a family can decrease the time needed to derive and recognize related words (Nagy et al., 1989). In Appendix H, we discuss in detail our treatments of words and potential issues. In summary, our Minnow models are somehow robust to these variations. Future work could take morphological features into consideration during training, and conduct targeted evaluations of morphological variations of the learned words.
Quantitative Evaluation of Generations
In the generative evaluation settings (Section 5.3 and 5.4), we used ROUGE-L, BERTScore F1, and LLM-as-a-Judge (GPT-4o in our case) for automatic quantitative evaluations. However, these evaluations are imperfect. For example, different metrics are suitable for generations in different lengths, and LLM evaluators are known to have biases and inconsistencies (Doostmohammadi et al., 2024; Stureborg et al., 2024). Future work should conduct careful human evaluations for futher validation to avoid these potential issues.
@@ Acknowledgements
We thank Najoung Kim for providing the syntactic category classification dataset. We thank Ryan Teehan for providing the code and checkpoints for the CoLLEGe model. We also thank Michael Hu, Will Merrill, Sophie Hao, Byung-Doh Oh, Shauli Ravfogel, and other members of the Computation and Psycholinguistics Lab for insightful and helpful discussions and comments. This work is supported by the National Science Foundation under NSF Award 1922658 (for Wentao Wang) and IIS-2239862. This work is also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
@@ Appendix A Word Usage Dataset Creation
Since the basic units of our focus are words (as opposed to word pieces in other tokenization schemes), we need to identify words in the text. To achieve this, we apply the same word-level tokenization to all datasets (for consistency) and mark word boundaries by whitespace during preprocessing. Models trained from scratch use this word-level tokenization. When the text is used in finetuning Llama-3, which comes with its pre-trained subword tokenizer, we remove the unnatural spaces introduced by the word-level tokenization and tokenize the text again with Llama-3 tokenizer, so the text format becomes closer to its pre-training data (See the Finetuning paragraph in Appendix B for further details of this process). For CHILDES data, we preprocess the data in the same way as Yedetore et al. (2023) did, which uses children’s input in the North American English portion, [^13] The version of CHILDES data we use is different from that of Yedetore et al. (2023), and the current version on the official webpage https://childes.talkbank.org/access/Eng-NA/ has also changed from our version. but we do not split and unk the data at the preprocessing stage. For BabyLM data, we use the data in the 10M track of the BabyLM Challenge 2023, which mixes 10 portions, each from a different data source (child- or adult-oriented, speech transcription or written text like Wikipedia). We exclude the QED portion for its poor quality (also mentioned in the 2nd BabyLM Challenge). We apply word-level tokenization on untokenized portions, and then split the text into sentences using heuristics. We use spaCy for all word-level tokenization along with Part-Of-Speech tagging. We lowercase all text before preprocessing to unify the capitalization of words in different places. We deduplicate sentences and remove sentences having less than 1 word (not counting punctuation).
@@ Appendix B Model and Training Configurations
We finetune Llama-3 8B (Meta AI, 2024) with Minnow on each of the CHILDES and BabyLM-10M datasets, but we refer to the models finetuned on BabyLM-10M by default, as we mentioned in Section 5. We finetune from both the pre-trained and instruction-tuned variants of Llama-3 8B, but we refer to the models finetuned from the pre-trained variant by default, presenting results of finetuning from the instruction-tuned variant only in the generative settings, where their performance may differ significantly due to their different capabilities to follow the prompt. We use two reserved special tokens in Llama-3 tokenizer vocabulary as the placeholder token and the separator token. To make the tokenization more natural to the model’s pre-training data, we clean up tokenization spaces in the text (e.g., the space before “,”, “.”, or “’s”) introduced by the word-level tokenization during preprocessing and make the placeholder token absorbs any preceding spaces of the word. Finetuning is minimally parameter-efficient: We finetune only the input and output embeddings of the two special tokens, while freezing all other parameters. Before finetuning, the input/output embedding of either token is initialized to the mean of all input/output embeddings (Hewitt, 2021). When finetuning the model on CHILDES with 5 examples per episode, we use batch size (i.e., number of episodes/sequences per batch) $32$ and initial learning rate $3\times 10^{-3}$ and truncate the sequence to the max length of $80$ tokens to control the memory usage. When finetuning the model on CHILDES with 10 examples per episode, we use batch size $8$ and initial learning rate $3\times 10^{-4}$ and truncate the sequence to the max length of $180$ tokens. When finetuning the model on BabyLM-10M with $5$ examples per episode, we use batch size $16$ and initial learning rate $1\times 10^{-3}$ and truncate the sequence to the max length of $160$ tokens. Other settings are the same as when training from scratch except that we do not apply weight decay. Each run is performed on a single A100 GPU for 15 epochs on CHILDES (33 hours) or 12 epochs on BabyLM-10M (48 hours).
@@ Appendix C Held-out Word Classification
An example task
Here we provide an example task to further explain the classification paradigm in Section 3.3. Assume $K=4,~{}C=2$, and we reuse the example words and sentences in Figure 1. The word “ski” has its three study examples concatenated into a sequence:
and a query example of the word “ski” is formatted as:
He will [new-word] past the pine trees. <sep>
The word “aardvark” has its three study examples concatenated into another sequence:
<sep> Look there’s an [new-word], it’s like an anteater <sep> See the [new-word] has a long snout for eating bugs. <sep> That must be the [new-word]’s house. <sep>
and a query example of the word “aardvark” is formatted as:
The [new-word] is hungry, it wants some snacks. <sep>
When classifying the word “ski”, we compare the conditional likelihood of its query example “He will [new-word] past the pine trees. <sep>” in the following two sequences:
<sep> Susie learned to [new-token] last winter <sep> People [new-token] on tall mountains where there’s lots of snow <sep> I saw Susie [new-token] fast down the snowy mountain <sep> He will [new-word] past the pine trees. <sep>
<sep> Look there’s an [new-word], it’s like an anteater <sep> See the [new-word] has a long snout for eating bugs. <sep> That must be the [new-word]’s house. <sep> He will [new-word] past the pine trees. <sep>
and we expect the conditional likelihood to be higher in the former sequence.
Task construction
@@
This is the baseline model introduced in Section 3.4.1. We follow the format described there and additionally prepend a prompt to make the performance better: “The following lines are lowercased example sentences using a new word ‘dax’ in random order, one per line:”. (We discuss the consequence of using a same pseudo-word in Appendix H.)
@@
dataset $K$ $C$ Minnow from scratch Llama-3 8B with ‘’ Llama-3 8B with ‘dax Llama-3 8B +Minnow Llama-2 7B with ‘dax Llama-2 7B +Minnow Llama-2 7B +CoLLEGe
CHILDES $5$ 4 72.3(1.6) 58.33 71.09 79.1(0.5) 70.06 79.4(0.3) 62.45
8 59.8(0.4) 46.49 60.02 70.4(0.2) 59.09 70.8(0.3) 47.93
$10$ 4 75.1(0.7) 66.56 76.53 84.9(0.2) 76.23 82.0(0.4) 63.80
8 63.4(1.5) 56.17 66.05 75.9(0.6) 65.74 73.1(0.4) 50.62
BabyLM-10M $5$ 4 77.4(0.5) 70.45 78.39 86.5(0.6) 78.34 85.8(0.5) 75.25
8 67.5(0.7) 60.12 69.74 80.5(1.0) 69.53 79.5(0.4) 63.56
@@ Appendix D Syntactic Category Classification
Figure 2: Syntactic classification accuracy. Error bar shows the 95% confidence interval given 3 runs. “Minnow from scratch on CHILDES” and “Minnow from scratch on BabyLM-10M” mean the models trained from scratch with Minnow on CHILDES and BabyLM-10M, respectively. (These models have a closed vocabulary, so many words in the dataset will be Out-Of-Vocabulary and be presented as <unk>, which could make the task easier.) “baseline” means baseline with pseudo-word “dax”, “wug”, or “blicket”. “+Minnow on BabyLM-10M” means Minnow finetuning on BabyLM-10M. “+CoLLEGe” means the CoLLEGe model (Teehan et al., 2024; generated new embeddings are used by Llama-2 7B). “N”, “V”, “Adj”, and “Adv” are short for noun, verb, adjective, and adverb, respectively. “Mean” is the mean across all category pairs. The black dashed line marks the chance level (50%). “Llama-3 8B +Minnow on BabyLM-10M” shows improvement over “Llama-3 8B baseline” in all category pairs, with mean accuracy risen from 66% to 83%. Meanwhile, both “Minnow” and “CoLLEGe” improve the accuracy of “Llama-2 7B baseline” from 69% to 80%. Note that “Minnow from scratch on BabyLM-10M” has a 77% mean accuracy, much better than the baseline accuracy and even comparable to the Minnow models finetuned from Llama-3 8B on many category pairs, again demonstrating its data efficiency.
@@
Noun Verb 78.7(4.4) 70.7(19.2) 86.7(12.0) 86.3(1.5) 74.7(1.7) 98.0(1.6)
Noun Adjective 65.2(2.1) 87.3(5.0) 43.0(1.6) 84.0(2.2) 71.3(4.6) 96.7(0.5)
Noun Adverb 61.0(2.9) 32.3(4.5) 89.7(4.0) 81.3(2.2) 75.7(1.7) 87.0(2.9)
Verb Adjective 74.2(5.2) 88.0(11.3) 60.3(18.1) 92.7(0.5) 90.0(2.2) 95.3(1.2)
Verb Adverb 56.3(0.6) 53.7(12.3) 59.0(11.9) 78.8(5.2) 90.0(2.4) 67.7(12.5)
Adjective Adverb 62.2(0.9) 60.3(10.4) 64.0(11.3) 72.8(0.2) 57.3(2.6) 88.3(3.1)
Table 7: LLMs’ accuracies (%) of distinguishing two syntactic categories in novel contexts. We show the mean and the standard deviation (in the bracket) of 3 runs. ‘Acc. (1$>$2)’ denotes the accuracy on the set of sentences where Category 1 should be preferred over Category 2 (e.g., assigning a higher probability to a noun in a noun-expecting context for row 1), and vice versa. Column ‘Acc.’ lists the aggregate accuracy. “Llama-3 8B +Minnow” have accuracies significantly better than chance except distinguishing adjective from verb (row 5). Additionally, “Llama-3 8B +Minnow” improves over Llama-3 8B baseline in differentiating most category pairs, showing the effectiveness of finetuning with Minnow.
@@
Table 8: Accuracies (%) of distinguishing two syntactic categories in novel contexts for models trained from scratch with Minnow. We show the mean and the standard deviation (in the bracket) of 3 runs. The formatting is the same as in Table 7. Both models perform better than chance on many category pairs, suggesting that models can develop some ability to one-shot learn the syntactic category of a word from human-scale data with Minnow.
Llama-2 7B +Minnow Llama-2 7B with CoLLEGe
Cat. 1 Cat. 2 Acc. Acc. (1$>$2) Acc. (2$>$1) Acc. Acc. (1$>$2) Acc. (2$>$1)
Noun Verb 84.7(1.6) 70.7(4.2) 98.7(0.9) 58.5 92 25
Noun Adjective 87.8(1.0) 80.3(2.5) 95.3(0.5) 91.0 84 98
Noun Adverb 75.3(0.6) 70.0(3.6) 80.7(2.6) 76.0 58 94
Verb Adjective 91.5(0.8) 95.0(1.4) 88.0(2.2) 94.5 99 90
Verb Adverb 78.0(1.6) 91.7(1.2) 64.3(2.1) 79.5 68 91
Adjective Adverb 60.2(1.0) 45.0(4.3) 75.3(4.9) 78.5 69 88
Table 9: CoLLEGe’s Teehan et al. (2024) accuracies (%) of distinguishing two syntactic categories in novel contexts. The formatting is the same as in Table 7.
@@ Appendix E Comparing Generations
Tables 1 and 10 show the results of comparing Llama-3 8B baseline (pre-trained to instruction-tuned) to the Minnow model finetuned from that baseline (with random seed 0) on new examples and definitions generated by greedy decoding and top-p=$0.92$, respectively.
@@
Llama-3 8B Instruct baseline 46 52 33